# Floquet-Lindblad Master Equation Approach to Open Quantum System Dynamics

### v1

@@ III.1 Driven Two-Level System
The first exemplar of the utility of FLiMESolve can be found in a relatively simple system: the driven 2-level system. Specifically considered here is a scenario wherein an oscillating electric or magnetic field is driving a general 2LS, such as a quantum dot or superconducting qubit. The Hamiltonian for this system can be written using quantum optics notation (and with $\hbar=1$) as [15, 16, 17, 18, 19, 20]:
$\hat{H}(t)=\frac{\omega_{0}}{2}\hat{\sigma}_{z}+\vec{E}(t)\cdot{}\hat{d}.$ (25)
In the above, $\omega_{0}$ is the transition frequency of the system, $\hat{\sigma}_{z}$ is the Pauli z operator, $\vec{E}(t)$, is the electric field of the driving laser, and $\hat{d}$ is the transition dipole moment of the 2LS.
@@
The Rabi frequency, $\Omega$, is the magnitude of the coupling between the forward-rotating laser frequency and the dipole moment. This transformation has the benefit of allowing one to rewrite the Hamiltonian in a time-independent and analytically solvable format in a rotating reference frame. Further, this formulation has a definite steady state in the computational basis, which is given by [23, 24]:
@@
However, while it is a very good approximation at optical frequencies where the laser coupling strength is usually much less than the resonance frequency, the RWA cannot be applied in systems where the driving strength of the field is an appreciable fraction of the resonance frequency of the system; i.e., where $\Omega\ll\omega_{0}$ is not true. When this is the case, the RWA is not valid, as the rapidly oscillating terms contribute substantially to the dynamics. The steady state predicted by the RWA becomes inaccurate, as the system instead approaches a nonequilibrium steady state, i.e., an oscillating state to which all initial states are attracted over time [12, 13].
@@
The new terms given by $\tilde{\Omega}$ are the coupling between the dipole moment and the “counter-rotating” components of the field. However, when the full system Hamiltonian is used, the equation of motion can no longer be analytically solved because of the time dependence of the Hamiltonian. The time dependence is periodic, however, which is handled efficiently by Floquet theory and FLiMESolve.
@@
Figure 1: State evolution of a driven 2LS. In the above, $\Delta=0$, $\gamma=2.5$ GHz. In (a), the laser driving strength ($\Omega/\omega_{0}=5\times 10^{-5}$) is low enough that a steady state is reached, which agrees well with the RWA steady-state population predicted by Eqn. 27. However, in (b), the laser driving power is half the resonance frequency of the 2LS ($\Omega/\omega_{0}=0.5$). In this case, the RWA steady state is no longer a good solution, and the full system dynamics must be accounted for.
@@
Further, because the system often relaxes on a timescale on the order of $10^{5}-10^{6}$ times longer than the period of the laser, the system must be simulated for an rather large number of periods to reach its nonequilibrium steady state (NESS). Figure 2 shows the evolution of the simulated 2LS when the Rabi frequency is equal to half the resonance frequency of the 2LS. Figure 2(a) shows the overall decay of the driven 2LS to a nonequilibrium steady state (NESS) over a few hundreds of thousands of periods of driving, while Fig. 2(b) shows the single-period dynamics that persist throughout the system evolution and into the NESS.
@@
Since the system must be simulated for so long to reach the limit cycle when straightforwardly solving the associated master equation, this results in a very large computational cost. However, as it is uniquely suited to simulating time-periodic systems, FLiMESolve can perform the simulation faster than, e.g., QuTiP’s MESolve. This is after some initial overhead investment to set up the internal rate matrices in FLiMESolve, perform basis changes to the Floquet basis, etc. Table 1 shows the computation time, in seconds, required for MESolve versus FLiMESolve for various simulation times. When the number of periods simulated is low, MESolve is slightly faster than FLiMESolve. As previously mentioned, this is due to the overhead required to set up FLiMESolve. However, as the number of periods simulated grows, FLiMESolve catches up to MESolve in computational time (at around 100 periods of simulated time), before surpassing MESolve in speed. FLiMESolve’s advantage appears to saturate at around 5.5-6.0 times faster than MESolve for large simulation times. This means that for time-periodic systems, FLiMESolve is either comparable or faster than MESolve for all simulation times. However, the exporting of these solutions requires applying QuTiP’s ”quantum object” wrapper around every state individually. This process reduces the speed of FLiMESolve majorly, causing the total run time to be roughly the same for both FLiMESolve and MESolve. However, FLiMESolve is currently written to inefficiently use Numpy for most of its data operations, and a rewriting of FLiMESolve in QuTiP’s data layer (which is a faster way to do data operations and which interacts better with the quantum object wrapper) should alleviate this bottleneck and restore the efficiency of FLiMEsolve over MESolve as seen in table 1. All simulations were performed using using an Intel i7-12700K processor and 16 GB RAM.
@@
Figure 2: Plots of the excited state evolution in time on two different timescales. (a) The system simulated for 500,000 periods of driving, with the nonequilibrium steady state occurring ar around 200,000 periods. (b) Two periods of oscillation in the nonequilibrium steady state, showing the single-period scale dynamics of the system.
@@
Number of Periods FLiMESolve Solution Time (s) MESolve Solution Time (s) Solution Time Quotient
$10^{1}$ $5.0\times 10^{-3}$ $2.0\times 10^{-3}$ 0.4
$10^{2}$ $7.0\times 10^{-3}$ $1.6\times 10^{-2}$ 0.89
$10^{3}$ 0.04 0.19 4.61
$10^{4}$ 0.30 1.71 5.43
$10^{5}$ 2.63 17.484 5.92
$10^{6}$ 28.87 181.87 5.62
@@
Table 1: Solution time and total run time in seconds for FLiMESolve and MESolve for the driven two-level system and varied numbers of periods of simulation. The rightmost column show the solution time of MESolve divided by the solution time of FLiMESolve (i.e., before applying the QObj() wrapper to the Floquet states and solution states)
III.2 $\pi$-Pulsed Two-Level System
@@
The next system of interest is again the 2-level system, however, this time, the excitation is a train of $\pi$-pulses. The effect of the pulses is to “invert” the populations of the ground and excited states, with the system evolving as a non-driven 2LS between pulses. To begin, consider the Hamiltonian given by
$\hat{H}(t)=\frac{1}{2}\begin{bmatrix}-\Delta&0\\ 0&\Delta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\sum_{j=0}^{\infty}\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(t-j*\epsilon)^{2}}{2\sigma^{2}}}$ (29)
@@
In the above equation, $\epsilon$ and $\sigma$ are defined from the lifetime of the system, defined by $\Gamma=2$ units of time (where here, the prefix is arbitrary due to the rest of the system time parameters being defined from $\Gamma$). Defining $T_{1}=\Gamma=2$, $\epsilon$ is chosen to be $0.05T_{1}$, and $\sigma$ is chosen to be $T_{1}/16$ This has the result of creating five pulses within any given lifetime of the system. The resulting pulse train can be seen in Fig. 3.
@@
Figure 3: The pulse train used to drive the 2LS. The magnitude of each pulse is meaningless besides the fact that each pulse is area-normalized to $\pi$. The pulse train is set up to pulse five times per lifetime of the system
III.2.1 Relaxed Secular Approximation
@@
The solved behavior of this system looks markedly different depending on the level of secular approximation used. Looking at Fig. 4 a, the state evolution is plotted with the most restrictive secular approximation.
@@
Figure 4: The state evolution as solved by FLiMESolve for the (left) the most restrictive secular approximation and (right) an effectively free secular approximation
@@
In this case, it appears that the overall “envelope” of the decay is present, however some nonphysical state evolution is also present. Rather than the excited state always decaying and the ground state always growing in population, the state with higher population is always decaying, and the state that has less population is always growing. This is clearly incorrect, as the system should act as a free 2LS between pulses, i.e., the excited state should always decay between pulses and the ground state should always grow between pulses. However, looking at the right side of Fig. 4, a more physically correct behavior can be seen. This figure shows the result of FLiMESolve handling the state evolution with an effectively infinite cutoff for the secular approximation. This figure shows much more physically accurate evolution, with the expected decay (growth) of the excited (ground) state population between pulses.

### v2

@@ III.1 Driven Two-Level System
The first exemplar of the utility of FLiMESolve can be found in a relatively simple system: the driven 2-level system. Specifically considered here is a scenario wherein an oscillating electric or magnetic field is driving a general 2LS, such as an atom, a quantum dot, or a superconducting qubit. The Hamiltonian for this system can be written using quantum optics notation (and with $\hbar=1$) as [16, 17, 18, 19, 20, 21]:
$\hat{H}(t)=\frac{\omega_{0}}{2}\hat{\sigma}_{z}+\vec{E}(t)\cdot{}\hat{\vec{d}}.$ (25)
In the above, $\omega_{0}$ is the transition frequency of the system, $\hat{\sigma}_{z}$ is the Pauli z operator, $\vec{E}(t)$, is the electric field of the driving laser, and $\hat{\vec{d}}$ is the transition dipole moment of the 2LS. In the following, we will assume that the electric field has the form
$\vec{E}(t)=\vec{E}\cos(\omega t+\phi)=\left(\vec{\epsilon}\mathcal{E}e^{i\omega t}+\vec{\epsilon}^{*}\mathcal{E}^{*}e^{-i\omega t}\right)$ (26)
where $\vec{\epsilon}$ is the (in general complex) polarization vector, and $\mathcal{E}$ is the complex electric field amplitude.
@@
The Rabi frequency, $\Omega=\mathcal{E}\vec{\epsilon}\cdot\vec{d}$, is the magnitude of the coupling between the forward-rotating laser frequency and the dipole moment. This transformation has the benefit of allowing one to rewrite the Hamiltonian in a time-independent and analytically solvable format in a rotating reference frame. Further, this formulation has a definite steady state in the computational basis, which is given by [24, 25]:
@@
However, while it is a very good approximation at optical frequencies where the laser coupling strength is usually much less than the resonance frequency, the RWA cannot be applied in systems where the driving strength of the field is an appreciable fraction of the resonance frequency of the system; i.e., where $\Omega\ll\omega$ is not true. When this is the case, the RWA is not valid, as the rapidly oscillating terms contribute substantially to the dynamics. The steady state predicted by the RWA becomes inaccurate, as the system instead approaches a nonequilibrium steady state, i.e., an oscillating state to which all initial states are attracted over time [13, 14].
@@
The new terms given by $\tilde{\Omega}=\mathcal{E}^{*}\vec{\epsilon}^{*}\cdot\vec{d}$ are the coupling between the dipole moment and the “counter-rotating” components of the field. However, when the full system Hamiltonian is used, the equation of motion can no longer be analytically solved because of the time dependence of the Hamiltonian. The time dependence is periodic, however, which is handled efficiently by Floquet theory and FLiMESolve.
@@
Figure 1: Time evolution of a driven 2LS with $\Delta=0$, and $\gamma=2.5$ GHz. (a) With a low laser-driving strength ($\Omega/\omega_{0}=5\times 10^{-5}$), a steady state is reached, which agrees well with the RWA steady-state population predicted by Eqn. (28). (b) With the Rabi frequency equal to half the resonance frequency of the 2LS ($\Omega/\omega_{0}=0.5$), the RWA steady state is no longer a good solution, and the full system dynamics must be accounted for.
@@
Further, because the system often relaxes on a timescale on the order of $10^{5}-10^{6}$ times longer than the period of the laser, the system must be simulated for a rather large number of periods to reach its nonequilibrium steady state (NESS). Figure 2 shows the evolution of the simulated 2LS when the Rabi frequency is equal to half the resonance frequency of the 2LS. Figure 2(a) shows the overall decay of the driven 2LS to a nonequilibrium steady state (NESS) over a few hundreds of thousands of periods of driving, while Fig. 2(b) shows the single-period dynamics that persist throughout the system evolution and into the NESS.
@@
Since the system must be simulated for so long to reach the nonequilibirum steady state when straightforwardly solving the associated master equation, this results in a very large computational cost. However, as it is uniquely suited to simulating time-periodic systems, FLiMESolve can perform the simulation faster than, e.g., QuTiP’s MESolve. This is after some initial overhead investment to solve for the Floquet modes and quasienergies, set up the internal rate matrices in FLiMESolve, perform basis changes to the Floquet basis, etc. Table 1 shows the solution time, in seconds, required for MESolve versus FLiMESolve for various simulation times. All simulations were performed using using an Intel i7-12700K processor and 16 GB RAM. When the number of periods simulated is low, MESolve is slightly faster than FLiMESolve. As previously mentioned, this is due to the overhead required to set up FLiMESolve. However, as the number of periods simulated grows, FLiMESolve catches up to MESolve in computational time (at around 100 periods of simulated time), before surpassing MESolve in speed. FLiMESolve’s advantage appears to saturate at around 5.5-6.0 times faster than MESolve for large simulation times. This means that for time-periodic systems, FLiMESolve is either comparable to or faster than MESolve for all simulation times. However, the exporting of these solutions requires applying QuTiP’s “quantum object” wrapper, QObj(), around the state individually at every time step. This process reduces the overall speed of FLiMESolve, causing the total run time to be only slightly shorter for FLiMESolve than MESolve. However, FLiMESolve is currently written using numpy for most of its data operations, and a rewriting of FLiMESolve in QuTiP’s data layer (which is a faster way to do data operations and which interacts better with the quantum object wrapper) should alleviate this bottleneck and restore the efficiency of FLiMEsolve over MESolve as seen in Table 1.
@@
Figure 2: Plots of the excited state evolution of the resonantly driven 2LS with $\Omega/\omega_{0}=0.5$. (a) The system simulated for 500,000 periods of driving, with the nonequilibrium steady state being reached at around 200,000 periods. (b) Two periods of oscillation in the nonequilibrium steady state, showing the single-period-scale dynamics of the system.
@@
Number of Periods FLiMESolve Solution Time (s) MESolve Solution Time (s) Solution Time Quotient Total Run Time Quotient
$10^{1}$ $5.0\pm 5.3\times 10^{-3}$ $2.0\pm 0.3\times 10^{-3}$ 0.4 0.39
$10^{2}$ $2.1\pm 0.2\times 10^{-2}$ $1.91\pm 3.3\times 10^{-1}$ 8.9 2.57
$10^{3}$ $0.04\pm 0.002$ $0.19\pm 0.008$ 4.61 1.14
$10^{4}$ $0.30\pm 0.012$ $1.71\pm 0.048$ 5.43 1.12
$10^{5}$ $2.63\pm 0.03$ $17.48\pm 0.48\pm$ 5.92 1.18
$10^{6}$ $28.87\pm 0.35$ $181.87\pm 2.27$ 5.62 1.23
@@
Table 1: Solution time and total run time in seconds for FLiMESolve and MESolve for the driven two-level system and varied numbers of periods of simulation. The rightmost two columns show the solution time of MESolve divided by the solution time of FLiMESolve (i.e., before applying the QObj() wrapper to the Floquet states and solution states) and the total run time of MESolve divided by the total run time of FLiMESolve (after applying the QObj wrapper to the solution and Floquet states), respectively.
III.2 Bichromatic Two-level System
The bichromatic 2LS is an interesting and dynamic system, and motivates the use of FLiMESolve due to the fact that even under the rotating wave approximation, the system cannot be made fully time-independent. As with the excited two level system, we first start with the Atomic Hamiltonian. Assuming that the normalized energy eigenstates of the system are such that $\hat{H}_{atom}\ket{\Psi_{1}}=E_{1}\ket{\Psi_{1}}$ and $\hat{H}_{atom}\ket{\Psi_{2}}=E_{2}\ket{\Psi_{2}}$, the atomic Hamiltonian can be written in the rotating wave approximation as
$\hat{H}_{atom}=\begin{bmatrix}-\Delta&0\\ 0&\Delta\end{bmatrix}$ (30)
, where $\Delta=E_{2}-E_{1}$ is the energy difference between the two levels of the system. Additionally, the same dipole operator is defined as before, namely that
[…]
Now it’s time to describe the atom-laser interaction. As with before, this is defined as the dot product between the dipole moment of the atom and the electric fields of the laser(s). However, unlike before, there are now two lasers for which to account:
$\hat{\vec{d}}\bullet\vec{E}(t)=\hat{\vec{d}}\bullet\vec{E}_{1}(t)+\hat{\vec{d}}\bullet\vec{E}_{2}(t)$ (32)
Now, the act of performing the rotating wave approximation amounts to ignoring the quickly rotating terms in Eqn. (32). In this case, that means ignoring terms with sums in the frequencies, i.e., only keeping terms that depend on the beat frequency. Performing the RWA and then separating the Hamiltonian based on time dependence, the overall Hamiltonian can be written as:
$\tilde{H}_{0}$ $=\frac{\bar{\Delta}}{2}\ \begin{bmatrix}-1&0\\ 0&1\end{bmatrix}$ (33)
$\tilde{H}^{+}(t)$ $=-\frac{1}{2}\ \begin{bmatrix}0&\Omega_{2}\\ \Omega_{1}^{*}&0\end{bmatrix}e^{\frac{ib}{2}t}$
$\tilde{H}^{-}(t)$ $=-\frac{1}{2}\ \begin{bmatrix}0&\Omega_{1}\\ \Omega_{2}^{*}&0\end{bmatrix}e^{-\frac{ib}{2}t}$
The discussion portion of the Bichromatic two level system will solely concern itself with the emission spectra of the system. This is to more clearly show the parallels between simulation results produced by FLiMESolve and the results produced by Chris Gustin et. al. [26]. The simulated system in this simulation will use the same numbers as Gustin’s experimental system, i.e., that the resonance of the two level system is at $\frac{\omega_{0}}{2\pi}=330$ THz, and that the lifetime of the system is $455$ ps.
In all simulations, $\Omega_{1}$ is set to be $\frac{\Omega_{1}}{2\pi}=30\mu$eV Fig. 3 a and b show singular emission spectra for two separate values of $\Omega_{2}$. The top spectrum in each plot is found from $\frac{\Omega_{2}}{2\pi}=40\;\mu$eV, and the bottom plot in both is found using $\frac{\Omega_{2}}{2\pi}=20\;\mu$eV. Additionally, in the FLiMESolve-produced plots, three vertical lines are added to show the locations of the Mollow triplet sidebands that would be produced by $\Omega_{1}$ alone [27].
On the $\Omega_{2}=20\;\mu$eV plots, the ”original” Mollow triplet bands produced by Laser 1 are still largely visible. However, the main peak of the Mollow triplet has diminished, and the right sideband peak is also small compared to some new peaks around the original Mollow triplet peak locations. These new peaks are created by the second laser, which is detuned exactly to the left Mollow triplet peak. This second laser causes each original Mollow triplet peak to split into its own triplet. This can be seen by the peaks to the left and the right of the main Mollow triplet peaks which are highlighted by the vertical lines in the FLiMESolve plots. Overall, the peak locations and relative sizes match up very well between FliMESolve and Gustin’s bespoke simulations for this specific setup.
Moving on to the upper half (or the $\Omega_{2}=40\;\mu$eV parts of each image), the ”new” triplet peaks begin to overtake the main Mollow triplet peaks. What at first appears to be a shifted triplet in Gustin’s figure, located at around $10\;\mu$eV on the x axis, reveals itself to actually be formed of the new sidebands from the ”main” Mollow triplet as viewed in the FLiMESolve plots. The relative peak sizes and the exact peak frequencies match up with Gustin’s results almost exactly, barring the peaks with the same spectral width as the laser. These peaks are likely laser scattering lines from the simulated quantum dot.
Figure 3: a A figure from [26] showing the emission two singular emission spectra for the bichromatic 2LS. The top spectrum uses $\Omega_{2}=40\;\mu$eV, while the bottom is found using $\Omega_{2}=20\;\mu$eV. The Orange line is the spectrum found using Gustin’s bespoke simulation, and the black line is experimental data found by Gustin. b The FLiMESolve simulated results of the bichromatic system using (top) $\Omega_{2}=40\;\mu$eV and (bottom) $\Omega_{2}=20\;\mu$eV.
Finally, the speed at which MESolve and FLiMESolve are able to simulate the system is once again compared.
Number of Periods FLiMESolve Solution Time (s) MESolve Solution Time (s) Solution Time Quotient Total Run Time Quotient
$10^{1}$ $1.2\pm 0.6\times 10^{-2}$ $6\pm 7\times 10^{-3}$ 2.41 0.49
$10^{2}$ $7\pm 11\times 10^{-2}$ $5\pm 2\times 10^{-2}$ 2.26 0.70
$10^{3}$ $1.5\pm 0.08\times 10^{-1}$ $0.49\pm 0.02\times 10^{-1}$ 3.24 0.78
$10^{4}$ $1.67\pm 0.02$ $5.04\pm 0.19$ 3.01 0.79
$10^{5}$ $16.47\pm 0.49$ $49.33\pm 1.33$ 2.99 0.79
Table 2: Solution time and total run time in seconds for FLiMESolve and MESolve for the bichromatic two level system.
III.3 $\pi$-Pulsed Two-Level System
@@
The final system of interest is again the 2-level system, however, this time, the excitation is a train of $\pi$-pulses. The effect of the pulses is to invert the populations of the ground and excited states, with the system evolving as a non-driven 2LS between pulses. To begin, consider the rotating wave approximation form of the Hamiltonian given by
$\hat{H}(t)=\frac{1}{2}\begin{bmatrix}-\Delta&0\\ 0&\Delta\end{bmatrix}+\frac{1}{2}\begin{bmatrix}0&1\\ 1&0\end{bmatrix}\sum_{n=0}^{\infty}\frac{1}{\sigma\sqrt{2\pi}}e^{-\frac{(t-nT)^{2}}{2\sigma^{2}}}$ (34)
@@
In the above equation, the period $T$ and the Gaussian standard deviation $\sigma$ are quantified relative to the lifetime of the system, defined by $T_{1}=2$ units of time, where the exact units don’t matter because the Hamiltonian is set up to be scale invariant. The period is chosen to be $T=T_{1}/20$, and the Gaussian width is chosen to be $\sigma=T_{1}/16$ This has the result of creating five pulses within any given lifetime of the system. The resulting pulse train can be seen in Fig. 4.
@@
Figure 4: The pulse train used to drive the 2LS. The magnitude of each pulse is meaningless besides the fact that each pulse is area-normalized to $\pi$. The pulse train is set up to pulse five times per lifetime of the system.
III.3.1 Relaxed Secular Approximation
@@
The solved behavior of this system looks markedly different depending on the level of secular approximation used. In Fig. 5(a), the state evolution is plotted with the most restrictive secular approximation.
@@
Figure 5: The state evolution as solved by FLiMESolve for (a) the most restrictive secular approximation and (b) effectively no secular approximation
@@
In this case, it appears that the overall “envelope” of the decay is correct, however some nonphysical state evolution is also present. Rather than the excited state always decaying and the ground state always growing in population between the excitation pulses, the state with higher population is always decaying, and the state that has less population is always growing. This is clearly incorrect, as the system should act as a free 2LS between pulses, i.e., the excited state should always decay between pulses and the ground state should always grow between pulses. However, in Fig. 5(b), a more physically correct behavior can be seen. This figure shows the result of FLiMESolve handling the state evolution with an effectively infinite cutoff for the secular approximation. This figure shows much more physically accurate evolution, with the expected decay (growth) of the excited (ground) state population between pulses.
Finally and once more, the solution times for this system are compared between MESolve and FLiMESolve. Contrary to the earlier examples, FLiMESolve is faster than MESolve for low numbers of simulation periods, but becomes slower as the number of simulated periods is increased. This is likely due to the fact that the secular approximation must be relaxed to achieve accurate results in the case of FLiMESolve. Because the resulting dynamics are dependent on many more frequencies than just the central frequency of the system, FliMESolve’s strengths can’t be leveraged as well and the solution method takes longer. However, the ability of FLiMESolve to relax to allow for more complicated time dependence is a noteworthy benefit to the solution method.
Number of Periods FLiMESolve Solution Time (s) MESolve Solution Time (s) Solution Time Quotient Total Run Time Quotient
$10^{1}$ $0.021\pm.0.01$ $0.15\pm 0.014$ 7.29 5.64
$10^{2}$ $0.12\pm 0.007$ $0.49\pm 0.015$ 3.79 2.77
$10^{3}$ $0.97\pm 0.04$ $0.71\pm 0.04$ 0.73 0.047
$10^{4}$ $9.02\pm 0.1$ $2.63\pm 0.07$ 0.29 0.19
$10^{5}$ $89.12\pm 0.16$ $21.27\pm 0.69$ 0.23 0.16
Table 3: Solution time and total run time in seconds for FLiMESolve and MESolve for the $\pi$-pulsed two level system.